OpenStudy (anonymous):

Calculate K for the oxidation go Iron by H+ (Have worked this problem several times, followed all the appropriate steps and am not getting the same answer as the back of the book)

OpenStudy (anonymous):

Answer is: 4.5 * 10^3

OpenStudy (anonymous):

I used the equation \[\log_{K} = Ecell \left( \frac{ n }{ 0.0592 V } \right)\] for n i substituted 6 & my Ecell value was = 0.036

OpenStudy (anonymous):

i think what you did is right just continue the problem.. log(k)=3.65 k=10^(3.65)=10^(.65)*10^3= 4.6*10^3